OpenStudy (anonymous):

135 = 9/2(x+28) solve for x Medal and follow its 9over 2

OpenStudy (anonymous):

Plz help

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Uhhhh

OpenStudy (campbell_st):

multiply both sides of the equation by 2 then divide both sides by 9 it should be obvious after that

OpenStudy (anonymous):

What

OpenStudy (anonymous):

30 and 6.2

OpenStudy (campbell_st):

135 = 9/2 (x + 28) multiply both sides of the equation by 2 to remove the denominator in the fraction 270 = 9(x + 28) now divide both sides of the equation by 9 30 = x + 28 hopefully this makes sense and the answer is obvious

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ohhhhhhhhhh thx